0.5(6x + 2) = 1.5 - (x + 2)
Answer by gwendolyn(128)

0.5(6x + 2) = 1.5 - (x + 2)
First, let's multiply both sides by 2 to get rid of the decimals:
6x + 2 = 3 - 2(x + 2)
Next, let's distribute the -2 to the inside of the parentheses:
6x+2=3-2x-4
6x+2=(-2x)-1
Then, let's add 2x to both sides:
8x+2=-1
After that, we can subtract 2 from both sides:
8x=-3
Finally, we divide both sides by 8:
8x/8 = (-3)/8
x=-3/8
So, x is equal to -3/8.